# Why They Link: An Intent Taxonomy for Including Hyperlinks in Social Posts

### v1

@@ 4.2. Study 1: Intention Distribution Analysis
Task Design, Participants, and Study Context. Inspired by prior work (Snow et al., 2008; Hettiachchi et al., 2022), we design clear task instructions and an intuitive annotation interface. We adopt an iterative task refinement process, progressively enriching the information provided to annotators—from tweet text with URLs, to tweet text with the title and textual content of the linked URLs, and finally to the complete tweet context (including hashtags and mentions) together with the title and textual content of the linked URLs. This iterative refinement substantially reduces the annotation error rate from 32.0% to 16.5%. In addition, we conduct multiple pilot studies to verify that participants fully understand the task requirements and can reliably assess the intentions underlying URL sharing in tweets. We require each tweet to be annotated by the same five eligible workers to simplify the inter-rater reliability calculations. To ensure that, we published the tweets in a sequence of four batches. Each batch was open for workers who completed the previous batch. The first batch had 100 tweets, and the workers were paid \$0.06 per tweet. We increased the rewards in each following batch by \$0.02 and the number of tweets by 100. As a result of this reward scheme, the entire set was annotated by the same workers. It also speed up the annotation process since the workers were motivated to earn a higher reward. The total cost of the annotation was about \$750 meaning that each worker received \$150 to annotate 1000 tweets.
@@
Results. On average, workers took 161.8 seconds to annotate a tweet with a standard deviation of 90.1 seconds. The average times of our five workers (in seconds) were 114.1, 125.4, 162.2, 175.3, and 226.9. Naturally, tweets vary in length and complexity. Hence, some workers needed more time to understand and annotate them, which explains the high standard deviation. The resulting annotation is shown in Figure 4. We label the intention of a URL in a tweet based on the high consensus. 22.5% of the tweets with URL have no high consensus (we discuss them in Sections 4.3). The most common intentions for tweets with URLs are promote, converse, share, and offer. Less than 1% of the tweets are annotated with request or entertain intentions. Finally, the workers give an uncertain intention for 4.4% of the tweets with URL. It is worth comparing that to 32.5%, which is the percentage of the tweets that receive the uncertain class using existing taxonomies from related work (Section 2.1). This shows that our taxonomy not only encompasses the known intentions identified in previous studies but also introduces novel categories that contribute to a more nuanced understanding of user intentions in the context of URL sharing.
@@ 4.3. Study 2: Effects of Expertise and Context on Annotation Behavior
The high consensus tweets in our result counts for 77.5%, including when the workers agree that the tweet is unclear (uncertain). The Fleiss’ kappa for all annotation items was $0.216$, and $0.259$ when only considering the high consensus tweets. The scores indicate fair agreement between annotators (Landis and Koch, 1977). In this study, we explore whether the expertise experience of annotators and more contextual information impact the annotation agreement.
@@ 6. Related Work
Communication researchers have developed a long list of why people communicate and much empirical work has been completed to assess the relative effectiveness of face-to-face, various forms of mediated-interpersonal (e.g., e-mail, phone), and a wide range of mass media channels concerning these functions. Flanagin and Metzger provides an exhaustive list of 21 reasons why people engage in various types of communication, and they find these 21 individual causes form nine clusters (Flanagin and Metzger, 2001). These clusters are as follows: information, learning, play, leisure, persuasion, social bonding, relationship maintenance, problem-solving, status, and insight. While the uses and gratifications research tradition has focused primarily on individuals as media consumers, it has been used to study why people use and what they get out of various social media experiences (e.g., (Chen, 2011; Phua et al., 2017; Smock et al., 2011)). Social media platforms allow individuals not only to be message consumers but also content creators. With this shift comes asking not just what information you get from social media sources, but what are the motivations behind and needs for creating social media content.
@@ 7. Future Work
Our future work includes a couple of directions: enhacing (microblog) information retrieval and integrating insights from communication science. In the former, we plan to extend the observations in Section 2.2 by embedding our taxonomy into a microblog retrieval system (Cambazoglu et al., 2021; Li et al., 2023), enabling intention-aware ranking and improved relevance for tasks such as credibility assessment and misinformation detection. In the latter direction, we aim to link our computational approach with theoretical perspectives from communication science, where communication intentions are central to understanding message effects. For example, research based on Reactance Theory investigates how audiences respond to perceived persuasive intent (Quick et al., 2015; Dillard and Shen, 2005), while studies on perceived partisan bias examine when audiences interpret news as persuasion rather than information (Eveland and Shah, 2003; Morris, 2007). These traditions highlight how intention distinguishes persuasion, social influence, and manipulation. Integrating such theoretical insights with computational modeling offers a promising interdisciplinary path for studying and modeling communicative intent in online discourse.
@@ 8. Conclusion
Understanding the intents of including URLs in posts provides an essential foundation for information retrieval on social media platforms, where URLs act as bridges connecting user-generated posts to the broader web. To systematically capture and model these purposes, we developed an intent taxonomy for including URLs in posts through a hybrid construction method that combines large-scale crowdsourced annotation with LLM assistance and conducted two studies, intention distribution and effects of expertise and context information for annotation agreement. Finally, we compare our intent taxonomy with those proposed in prior research and demonstrate a potential application. In a word, our work offers a reproducible model for building empirically grounded, LLM-augmented taxonomies applicable to future IR and NLP research.

### v2

@@ 4.2. Study 1: Intention Distribution Analysis
Task Design, Participants, and Study Context. Inspired by prior work (Snow et al., 2008; Hettiachchi et al., 2022), we design clear task instructions and an intuitive annotation interface. We adopt an iterative task refinement process, progressively enriching the information provided to annotators—from tweet text with URLs, to tweet text with the title and textual content of the linked URLs, and finally to the complete tweet context (including hashtags and mentions) together with the title and textual content of the linked URLs. This iterative refinement substantially reduces the annotation error rate from 32.0% to 16.5%. In addition, we conduct multiple pilot studies to verify that participants fully understand the task requirements and can reliably assess the intentions underlying URL sharing in tweets. We require each tweet to be annotated by the same five eligible workers to simplify the inter-rater reliability calculations. To ensure that, we published the tweets in a sequence of four batches. Each batch was open for workers who completed the previous batch. The first batch had 100 tweets, and the workers were paid \$0.06 per tweet. We increased the rewards in each following batch by \$0.02 and the number of tweets by 100. As a result of this reward scheme, the entire set was annotated by the same workers. It also sped up the annotation process since the workers were motivated to earn a higher reward. The total cost of the annotation was about \$750 meaning that each worker received \$150 to annotate 1000 tweets.
@@
Results. On average, workers took 161.8 seconds to annotate a tweet with a standard deviation of 90.1 seconds. The average times of our five workers (in seconds) were 114.1, 125.4, 162.2, 175.3, and 226.9. Naturally, tweets vary in length and complexity. Hence, some workers needed more time to understand and annotate them, which explains the high standard deviation. The resulting annotation is shown in Figure 4. We label the intention of a URL in a tweet based on the high consensus. 22.5% of the tweets with URL have no high consensus (we discuss them in Sections 4.3). The most common intentions for tweets with URLs are promote, converse, share, and offer. Less than 1% of the tweets are annotated with request or entertain intentions. Finally, the workers give an uncertain label for 4.4% of the tweets with URL. It is worth comparing that to 32.5%, which is the percentage of the tweets that receive the uncertain class using existing taxonomies from related work (Section 2.1). This shows that our taxonomy not only encompasses the known intentions identified in previous studies but also introduces novel categories that contribute to a more nuanced understanding of user intentions in the context of URL sharing.
@@ 4.3. Study 2: Effects of Expertise and Context on Annotation Behavior
The high consensus tweets account for 77.5% of our results, including when the workers agree that the tweet is unclear (uncertain). The Fleiss’ kappa for all annotation items was $0.216$, and $0.259$ when only considering the high consensus tweets. The scores indicate fair agreement between annotators (Landis and Koch, 1977). In this study, we explore whether the expertise experience of annotators and more contextual information impact the annotation agreement.
@@ 6. Related Work
Communication researchers have developed a long list of why people communicate and much empirical work has been completed to assess the relative effectiveness of face-to-face, various forms of mediated-interpersonal (e.g., e-mail, phone), and a wide range of mass media channels concerning these functions. Flanagin and Metzger provide an exhaustive list of 21 reasons why people engage in various types of communication, and they find these 21 individual causes form nine clusters (Flanagin and Metzger, 2001). These clusters are as follows: information, learning, play, leisure, persuasion, social bonding, relationship maintenance, problem-solving, status, and insight. While the uses and gratifications research tradition has focused primarily on individuals as media consumers, it has been used to study why people use and what they get out of various social media experiences (e.g., (Chen, 2011; Phua et al., 2017; Smock et al., 2011)). Social media platforms allow individuals not only to be message consumers but also content creators. With this shift comes asking not just what information you get from social media sources, but what are the motivations behind and needs for creating social media content.
@@ 7. Future Work
Our future work includes a couple of directions: enhancing (microblog) information retrieval and integrating insights from communication science. In the former, we plan to extend the observations in Section 2.2 by embedding our taxonomy into a microblog retrieval system (Cambazoglu et al., 2021; Li et al., 2023), enabling intention-aware ranking and improved relevance for tasks such as credibility assessment and misinformation detection. In the latter direction, we aim to link our computational approach with theoretical perspectives from communication science, where communication intentions are central to understanding message effects. For example, research based on Reactance Theory investigates how audiences respond to perceived persuasive intent (Quick et al., 2015; Dillard and Shen, 2005), while studies on perceived partisan bias examine when audiences interpret news as persuasion rather than information (Eveland and Shah, 2003; Morris, 2007). These traditions highlight how intention distinguishes persuasion, social influence, and manipulation. Integrating such theoretical insights with computational modeling offers a promising interdisciplinary path for studying and modeling communicative intent in online discourse (Alshehri et al., 2021, 2023).
@@ 8. Conclusion
Understanding the intents of including URLs in posts provides an essential foundation for information retrieval on social media platforms, where URLs act as bridges connecting user-generated posts to the broader web. To systematically capture and model these purposes, we developed an intent taxonomy for including URLs in posts through a hybrid construction method that combines large-scale crowdsourced annotation with LLM assistance and conducted two studies examining intention distribution and the effects of annotator expertise and contextual information on annotation agreement. Finally, we compare our intent taxonomy with those proposed in prior research and demonstrate a potential application. Overall, our work offers a reproducible model for building empirically grounded, LLM-augmented taxonomies applicable to future IR and NLP research.
Acknowledgements.
This work was supported in part by the U.S. National Science Foundation awards III-2107213 and 2026513.